OpenStudy (anonymous):

Find an equation of the sphere with center (3, -1, -1) and radius 4

myininaya (myininaya):

\[(x-h)^2+(y-k)^2+(z-l)^2=r^2\]

myininaya (myininaya):

where the center is \[(h,k,l)\] and r is radius

myininaya (myininaya):

just fill in the blanks: \[(x-\_)^2+(y-\_)^2+(z-\_)^2=\_^2\]

OpenStudy (anonymous):

(x-3)^2+(y+1)^2+(z+1)^2 =16

OpenStudy (anonymous):

\[(x-3)^2+(y+1)^2+(z+1)^2=16\]